A ball of mass 8kg falls from a height of 100m . Neglecting air resistance, calculate its total energy after falling a distance of 40m

1
Expert's answer
2022-03-24T11:01:24-0400

The total energy of a ball

"E=E_k+E_p=const"

"E=E_{k1}+E_{p1}=0+mgh_1\\
=8*9.8*100=7840\:\rm J"